Parth (parthkohli):

OpenStudy user L is 5 years older than OpenStudy user P. In 30 years, OpenStudy user P would be 6 more years than 3 times his current age. What would be the age of OpenStudy user L after 50 years from now?

OpenStudy (anonymous):

Lol you're still doing these? xD

Parth (parthkohli):

lol I couldn't get over them :P

OpenStudy (anonymous):

L=5+P P+30=6+3P 2P=24 P=12 L=5+P=17 L+50=67

OpenStudy (anonymous):

The answer is 67.

Parth (parthkohli):

Yes.

Parth (parthkohli):

CORRECT-O!

OpenStudy (anonymous):

But we all knew that.

Parth (parthkohli):

hha